An unfair coin has a probability 𝑝 = 0.13 of landing on heads when flipped. The coin is flipped 𝑛 = 8 times, and the proportion of heads is recorded. If this experiment is repeated many times, what are the mean 𝜇 and standard error 𝜎 of the distribution of the sample proportions?

Answer by stanbon(75887)

An unfair coin has a probability 𝑝 = 0.13 of landing on heads when flipped. The coin is flipped 𝑛 = 8 times, and the proportion of heads is recorded. If this experiment is repeated many times, what are the mean 𝜇 and standard error 𝜎 of the distribution of the sample proportions?
----
E(# of heads) = n*p = 8*0.13 = 1.04
std = sqrt(n*p*q) = sqrt(1.04*0.87) = 0.9512
-------------
Cheers,
Stan H.
-----------